Solving systems of equations using matrices:
4x + y = 7
-5x - y + z - w = 1
z+w = 7
I have z = 7 and know that w is any real number so far but am stuck solving for x and y. Help please.

Answer by Edwin McCravy(20086)


Write the equations this way

 4x +  y         = 7
-5x -  y + z - w = 1
           z + w = 7 

Then put 1's in front of the letters
without coefficients showing, and +0
times the missing letters:

 4x + 1y + 0z + 0w = 7
-5x - 1y + 1z - 1w = 1
 0x + 0y + 1z + 1w = 7 

Erase all the letters and
+ signs, move the minus
signs next to the coefficients
as negative signs, draw a line
where the equal signs are, and
put brackets around:

[ 4  1  0  0 | 7]
[-5 -1  1 -1 | 1]
[ 0  0  1  1 | 7]

Get a 0 where the -5 is by
adding 5 times the 1st row to
4 times the second row, and
restore the 1st row

5[ 4  1  0  0 | 7]
4[-5 -1  1 -1 | 1]
 [ 0  0  1  1 | 7]

 [ 4  1  0  0 |  7] 
 [ 0  1  4 -4 | 39]
 [ 0  0  1  1 |  7]

Get a 0 where the 1 in row 
1 is by adding -1 times row
2 to 1 times row 1, and
restore the 2nd row


 1 [ 4  1  0  0 |  7] 
-1 [ 0  1  4 -4 | 39]
   [ 0  0  1  1 |  7] 

   [ 4  0 -4  4 |-32] 
   [ 0  1  4 -4 | 39]
   [ 0  0  1  1 |  7] 

Get a 0 where the -4 is in
the 1st row by multiplying
the 3rd row by 4 and adding
it to 1 times the 1st row,
and restore the 3rd row


  1[ 4  0 -4  4 |-32] 
   [ 0  1  4 -4 | 39]
  4[ 0  0  1  1 |  7]

   [ 4  0  0  8 | -4] 
   [ 0  1  4 -4 | 39]
   [ 0  0  1  1 |  7]

Get a 0 where the 4 is in 
the 2nd row by multiplying
the 3rd row by -4 and
adding it to 1 times the 
2nd row, and restore the 
3rd row:

   [ 4  0  0  8 | -4] 
  1[ 0  1  4 -4 | 39]
 -4[ 0  0  1  1 |  7]

   [ 4  0  0  8 | -4] 
   [ 0  1  0 -8 | 11]
   [ 0  0  1  1 |  7]

Get a 1 where the first 4 is
in the 1st row by multiplying
the 1st row through by 1/4

   [ 1  0  0  2 | -1] 
   [ 0  1  0 -8 | 11]
   [ 0  0  1  1 |  7]

This is in row reduced
echelon form.

Put the letters, plus signs,
and equal signs back in:

   1x + 0y + 0z + 2w = -1
   0x + 1y + 0z - 8w = 11
   0x + 0y + 1z + 1w =  7

Simplify to

    x           + 2w = -1
         y      - 8w = 11
              z +  w =    7 

Eliminate the spaces (close it up:

x + 2w = -1
y - 8w = 11
z +  w =  7

Solve for x, y and z

     x = -1 - 2w
     y = 11 + 8w
     z = 7 - w

w can be any number, say c, then

     x = -1 - 2c
     y = 11 + 8c
     z = 7 - c
     w = c

So the solution is

(x, y, z, w) = (-1-c, 11+8c, 7-c, c) 

It does turn out that w can be any number,
but as you see, z is not necessarily 7, 
unless c happens to be chosen to be 0.

Edwin